OpenStudy (anonymous):

A pizza shop has 300 pounds (4800 ounces of dough. A small pizza uses 12 ounces of dough and a large pizza uses 18 ounces of dough. Write and graph and inequality describing the possible numbers of small and large pizzas that can be made. Then five three possible solutions.

OpenStudy (anonymous):

@jim_thompson5910 could you help again please/

jimthompson5910 (jim_thompson5910):

What are your variables here?

jimthompson5910 (jim_thompson5910):

ie, what are your unknowns?

OpenStudy (anonymous):

is it something like 12<x<18?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

Let x = number of small pizzas and y = number of large pizzas

OpenStudy (anonymous):

18=a(12)?

jimthompson5910 (jim_thompson5910):

no this is a completely different problem than before

OpenStudy (anonymous):

is it somtething like ethis? y-18=m(x-12)

jimthompson5910 (jim_thompson5910):

no it's simply this if you make x small pizzas and they take 12 oz of dough, then in total, the small pizzas will take 12x oz of dough

jimthompson5910 (jim_thompson5910):

use similar logic to find how much dough is needed for the large pizzas

OpenStudy (anonymous):

i dont get it..

OpenStudy (anonymous):

sorry

jimthompson5910 (jim_thompson5910):

focus on the small pizzas only for a moment

jimthompson5910 (jim_thompson5910):

1 sm pizza needs 12 oz 2 sm pizzas need 12*2 oz 3 sm pizzas need 12*3 oz ...etc etc x small pizzas need 12*x or just 12x ounces of dough making (more) sense now?

OpenStudy (anonymous):

12* x *18<300 is it somethinglike this?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

you need 18y ounces of dough to make 'y' large pizzas

jimthompson5910 (jim_thompson5910):

in total, you need 12x+18y ounces of dough to make x smalls and y larges

jimthompson5910 (jim_thompson5910):

you have 4800 oz of dough total

jimthompson5910 (jim_thompson5910):

so 12x+18y <= 4800 6(2x+3y) <= 4800 2x+3y <= 4800/6 2x+3y <= 800

OpenStudy (anonymous):

wait so the first part, inequality is 12x+18y <4800?

jimthompson5910 (jim_thompson5910):

12x+18y <= 4800

jimthompson5910 (jim_thompson5910):

<= means "less than or equal to"

jimthompson5910 (jim_thompson5910):

and 12x+18y <= 4800 simplifies to 2x+3y <= 800

OpenStudy (anonymous):

ok so do i do it like this? 2(12)+3(18)<800 24+54 78<800

jimthompson5910 (jim_thompson5910):

you just plugged in x = 12 and y = 18 this means that you're making 12 smalls and 18 larges

jimthompson5910 (jim_thompson5910):

since 78 is less than 800, this means that you have not used up all the available dough

OpenStudy (anonymous):

so what should i do?

jimthompson5910 (jim_thompson5910):

you need to graph

OpenStudy (anonymous):

could you help my graph? :/ so sorry

jimthompson5910 (jim_thompson5910):

graph 2x + 3y = 800 then shade below the line

OpenStudy (anonymous):

you mean graph y=-2/3+800/3?

jimthompson5910 (jim_thompson5910):

or you could graph y=-2/3x+800/3 yes